OpenStudy (anonymous):

Determine the period of the sin function y=0.75cos pi/2 x-0.5 a)4 b)pi/2 c)3/4 d)0.5

OpenStudy (phi):

\[ y= A cos(\frac{2 \pi}{T} x +\phi)\] where A is the amplitude T is the period \(\phi\) is the phase to find the period, equate 2pi/T to the coefficient of x inside the cos \[ \frac{2 \pi}{T}= \frac{\pi}{ 2} \] and solve for T

OpenStudy (anonymous):

do i cross multiply